# Wasserstein Evolution : Evolutionary Optimization as Phase Transition

### v1

@@ 5  Conclusion
This paper has established a novel theoretical and algorithmic framework that bridges evolutionary optimization with statistical physics through the lens of phase transition theory. We have demonstrated that the convergence dynamics of evolutionary algorithms can be rigorously formalized as a free energy minimization process via Wasserstein gradient flows. Building upon this foundation, we proposed Wasserstein Evolution (WE), an algorithm that explicitly models the exploration-exploitation trade-off as a competition between potential gradient forces and entropic forces, mirroring the disorder-to-order transition observed in physical systems.
The core contribution of this work is threefold. First, we established a formal mathematical correspondence between evolutionary optimization and statistical phase transitions, providing a unified theoretical perspective on population-based search algorithms. Second, we designed WE, an algorithm that implements the Wasserstein gradient flow of free energy, naturally balancing exploration and exploitation through temperature-controlled entropy dynamics. Third, extensive experiments on classical benchmark functions validated WE’s superior performance, particularly its ability to maintain significantly higher population diversity while achieving competitive or better convergence compared to established algorithms like GA, DE, and CMA-ES.

### v2

@@ 5  Conclusion
This paper establishes a novel theoretical and algorithmic framework that bridges evolutionary optimization with statistical physics through the lens of phase transition theory. We rigorously formalize the convergence dynamics of evolutionary algorithms as a free energy minimization process via Wasserstein gradient flows. Building upon this foundation, we propose Wasserstein Evolution (WE), an algorithm that explicitly models the exploration-exploitation trade-off as a competition between potential gradient forces and entropic forces, mirroring the disorder-to-order transition observed in physical systems. Theoretical analysis demonstrates that WE converges to the Boltzmann distribution, ensuring a principled balance between solution quality and diversity.
The core contributions of this work are threefold. First, we establish a formal mathematical correspondence between evolutionary optimization and statistical phase transitions, providing a unified theoretical perspective on population-based search. Second, we design WE, which implements the Wasserstein gradient flow of free energy, naturally balancing exploration and exploitation through temperature-controlled entropy dynamics. Third, extensive experiments validate WE’s superior performance against five established algorithms, including GA, DE, CMA-ES, JADE, and SaDE, on a diverse suite of benchmark and physical potential functions. The results show that WE consistently achieves the lowest free energy and maintains the highest entropy upon convergence. Furthermore, systematic invariance tests on transformed variants of the Schwefel 2.22 function confirm that WE possesses translation, scale, and rotation invariance, demonstrating its robustness and intrinsic alignment with the problem geometry rather than its specific parameterization.